### v2

@@ Abstract
Robots increasingly fulfill crucial roles in various tasks, such as planetary exploration and complex manipulation [1, 2, 3, 4]. Quadrupedal robots offer unparalleled flexibility and the capability to navigate challenging terrains. Fundamental to their interaction with the real world are advanced locomotion and navigation systems. In terms of locomotion, Model Predictive Control (MPC)-based methods [5] model environmental dynamics to achieve precise control. Reinforcement Learning (RL)-based methods [6, 7, 8, 9], trained in simulated environments like Isaac Gym [10], utilize curriculum learning [11] and domain randomization to enhance robustness and versatility. However, most locomotion policies that track velocity commands face significant tracking errors, which undermine their suitability for navigation tasks. Existing navigation approaches often use a hierarchical architecture where a high-level planner generates guidance signals based on environmental information like maps [12, 13, 14, 15, 16]. While some studies [17, 18, 19, 20] have incorporated robot feedback to aid navigation, the integration of a versatile locomotion policy with a general planner remains largely unexplored.
@@ 2.1 Quadrupedal Locomotion
Classical quadrupedal locomotion methods typically rely on Model Predictive Control (MPC), which demands high-precision dynamic modeling [5, 27]. Alternatively, RL-based methods have shown remarkable performance in traversing complex terrains using simulated learning and proprioceptive information [8, 9, 28, 29, 30, 31, 32, 33]. The integration of external sensors such as cameras and LiDAR with RL-based methods has further enhanced performance [34, 35, 36, 37]. Additionally, hybrid approaches that combine the strengths of MPC and RL have improved both accuracy and generalization of locomotion [38, 39]. Advanced maneuvers, including robot parkour, are now possible through RL and carefully designed reward functions [21, 22].
@@ 2.2 Hierarchical Navigation Architecture
Our goal is to integrate our versatile locomotion policy with general planners to navigate complex terrains effectively, without a complicated training process. Recent studies suggest that both classical planning algorithms and large language models (LLMs) can serve as effective planners, inspiring us to use them to generate waypoints and guide our locomotion policy [51, 52, 50].
@@ 3.2 Waypoint-guided Locomotion Policy
Fig. 2 illustrates the entire training procedure. The teacher policy is initially trained in a fixed waypoints (WP-Fixed) scenario and subsequently fine-tuned in a random waypoints (WP-Random) scenario. In contrast, the student policy is exclusively trained in the WP-Random scenario. Our policy outperforms those trained solely in one of the aforementioned scenarios, as demonstrated in Fig. 4.3.
@@ 4.1 Simulation Setup
We utilize Isaac Gym to train our locomotion policy. Specifically, for the WP-Fixed scenario, we create 10 rows and 40 columns of tracks, where each track contains 6 terrain units of the same type; and the WP-Random scenario consists of 10 rows and 10 columns of areas, where each area containing 6×5 terrain units. Table 1 provides the parameter ranges for critical obstacle properties such as box height and gap width. In the WP-Random scenario, we apply constraints on the generation of waypoints and the arrangement of terrain units to ensure that each waypoint remains accessible to the robot. For sim-to-real transfer, we incorporate domain randomization and observation noise throughout the training process, similar to approaches used in other open-source works [21, 9]. We train the teacher policy for 4000 iterations in the WP-Fixed scenario and 8000 iterations in the WP-Random scenario, followed by policy distillation to train the student policy for 6000 iterations. All training is conducted on an RTX-4090 GPU. We adopt Unitree AlienGo for simulation and real-world experiments.
@@ 4.2 Hardware Deployment
For real-world experiments, we deploy the Unitree Aliengo, equipped with a Jetson Orin NX for onboard computation and an Intel RealSense D435 for capturing depth images. We enhance and denoise images to align the depth data from simulations with real-world observations. The robot features 12 degrees of freedom. Our control policy operates at 50Hz and utilizes a built-in PD controller at 200Hz with with Kp = 40 and Kd = 2.0 for robot control. Additionally, a motion capture system is used to provide localization and ensure stable deployment performance.
@@ 4.3 Locomotion Comparison
We conducted experiments with and without high obstacles for both test tasks. Results presented in Table 2 reveal that RMA is unable to traverse the terrain, and robust walking without locomotion skills is insufficient to address these tasks. In the single-traverse task without high obstacles, all methods performed well; Extreme Parkour completed the task the fastest due to specific training in similar terrains, despite suffering from significant tracking errors. The failure of Ours-s1 can be attributed to the limited and regular distribution of waypoints provided during training. Conversely, Ours-s2 fails because training directly in the WP-Random scenario led to a terrible gait characterized by excessive jumping, rendering it unsuitable for deployment. Additionally, Ours-s2 struggles to bypass wide high obstacles directly in its path. Despite the presence of high obstacles, our method outperforms the others in the omni-traverse task. This superiority is due to the policy being trained in both scenarios, enabling it to handle more irregular local navigation situations. In some instances, Extreme Parkour deviates from the commands and approaches obstacles that could have been avoided. Heatmaps in Fig. 5 illustrate the position visiting frequencies of different methods in the omni-traverse task. Both Ours-s1 and Extreme Parkour struggle to navigate irregular obstacles, whereas our method visits farther positions more frequently, demonstrating more powerful locomotion skills.
@@ 4.4 Navigation System
Fig. 6 showcases sequential snapshots of our system navigating in a simulation using GPT-4, and in a real-world scenario using the A* algorithm. In the simulation, the navigation terrain consists of multiple terrain units, each 2m 2m, featuring varied heights and obstacles. The robot is initialized in one terrain unit, with another designated as the navigation target. The LLM (specifically, GPT-4) receives prompts as described in 3.3.2 and outputs waypoints in the form of terrain unit indices. These waypoints, positioned at the centers of the corresponding terrain units, guide our low-level controller to reach the target while handle obstacles not accounted for by the LLM. However, this waypoint configuration often overlooks intricate terrain details, potentially leading to impractical waypoint placements, such as within high obstacles. Our low-level controller can struggle to recover from some of these bad cases such as a waypoint in the middle of the gap or at the edge of a box, even though it has not been trained for such anomalies.
@@ 5 Conclusion
In this work, we introduced the use of waypoints as an interface to connect a versatile quadrupedal robot with a high-level planner for navigation. The locomotion policy trained in two simulated scenarios is able to traverse complex terrains while avoiding obstacles, and the high-level planner such as LLMs enhances generalization across various environments. Experiments demonstrate that the proposed navigation architecture is able to complete navigation tasks in complex scenarios. Future works include developing an end-to-end policy to conduct locomotion and navigation simultaneously.

### v3

@@ Abstract
Artificial Intelligence (AI) algorithms, represented by deep learning, have made outstanding advances in various domains like computer vision [1, 2, 3, 4, 5]. Robots with AI also increasingly fulfill crucial roles in various tasks, such as planetary exploration and complex manipulation [6, 7, 8, 9, 10, 11]. Quadrupedal robots offer unparalleled flexibility and the capability to navigate challenging terrains. Fundamental to their interaction with the real world are advanced locomotion and navigation systems. In terms of locomotion, Model Predictive Control (MPC)-based methods [12] model environmental dynamics to achieve precise control. Reinforcement Learning (RL)-based methods [13, 14, 15, 16], trained in simulated environments like Isaac Gym [17], utilize curriculum learning [18] and domain randomization to enhance robustness and versatility. However, most locomotion policies that track velocity commands face significant tracking errors, which undermine their suitability for navigation tasks. Existing navigation approaches often use a hierarchical architecture where a high-level planner generates guidance signals based on environmental information like maps [19, 20, 21, 22, 23]. While some studies [24, 25, 26, 27] have incorporated robot feedback to aid navigation, the integration of a versatile locomotion policy with a general planner remains largely unexplored.
@@ 2.1 Quadrupedal Locomotion
Classical quadrupedal locomotion methods typically rely on MPC, which demands high-precision dynamic modeling [12, 34]. Alternatively, RL-based methods have shown remarkable performance in traversing complex terrains using simulated learning and proprioceptive information [15, 16, 35, 36, 37, 38, 39, 40]. The integration of external sensors such as cameras and LiDAR with RL-based methods has further enhanced performance [41, 42, 43, 44]. Additionally, hybrid approaches that combine the strengths of MPC and RL have improved both accuracy and generalization of locomotion [45, 46]. Advanced maneuvers, including robot parkour, are now possible through RL and carefully designed reward functions [28, 29].
@@ 2.2 Hierarchical Navigation Architecture
Our goal is to integrate our versatile locomotion policy with general planners to navigate complex terrains effectively, without a complicated training process. Recent studies suggest that both classical planning algorithms and LLMs can serve as effective planners, inspiring us to use them to generate waypoints and guide our locomotion policy [58, 59, 57].
@@ 3.2 Waypoint-guided Locomotion Policy
Fig. 2 illustrates the entire training procedure. The teacher policy is initially trained in a fixed waypoints (WP-Fixed) scenario and subsequently fine-tuned in a random waypoints (WP-Random) scenario. In contrast, the student policy is exclusively trained in the WP-Random scenario. Our policy outperforms those trained solely in one of the aforementioned scenarios, as demonstrated in Sec. 4.3.
@@ 4.1 Simulation Setup
We utilize Isaac Gym to train our locomotion policy. Specifically, for the WP-Fixed scenario, we create 10 rows and 40 columns of tracks, where each track contains 6 terrain units of the same type; and the WP-Random scenario consists of 10 rows and 10 columns of areas, where each area containing 6×5 terrain units. Table 1 provides the parameter ranges for critical obstacle properties such as box height and gap width. In the WP-Random scenario, we apply constraints on the generation of waypoints and the arrangement of terrain units to ensure that each waypoint remains accessible to the robot. Notably, introducing noise during training plays a crucial role in enhancing the generalization ability of the policy [62, 63, 64, 65, 66]. For sim-to-real transfer, we incorporate domain randomization and observation noise throughout the training process, similar to approaches used in other open-source works [28, 16]. We train the teacher policy for 4000 iterations in the WP-Fixed scenario and 8000 iterations in the WP-Random scenario, followed by policy distillation to train the student policy for 6000 iterations. All training is conducted on an RTX-4090 GPU. We adopt Unitree AlienGo for simulation and real-world experiments.
@@ 4.2 Hardware Deployment
For real-world experiments, we deploy the Unitree Aliengo, equipped with a Jetson Orin NX for onboard computation and an Intel RealSense D435 for capturing depth images. We enhance and denoise images to align the depth data from simulations with real-world observations basically following [28, 29]. The robot features 12 degrees of freedom. Our control policy operates at 50Hz and utilizes a built-in PD controller at 200Hz with Kp = 40 and Kd = 2.0 for robot control. Additionally, a motion capture system is used to provide localization and ensure stable deployment performance.
@@ 4.3 Locomotion Comparison
We conducted experiments with and without high obstacles for both test tasks. Results presented in Table 2 reveal that RMA is unable to traverse the terrain, and robust walking without locomotion skills is insufficient to address these tasks. In the single-traverse task without high obstacles, all methods performed well; Extreme Parkour completed the task the fastest due to specific training in similar terrains, despite suffering from significant tracking errors. The failure of Ours-s1 can be attributed to the limited and regular distribution of waypoints provided during training. Conversely, Ours-s2 fails because training directly in the WP-Random scenario led to a terrible gait characterized by excessive jumping, rendering it unsuitable for deployment. Additionally, Ours-s2 struggles to bypass wide high obstacles directly in its path. Despite the presence of high obstacles, our method outperforms the others in the omni-traverse task. This superiority is due to the policy being trained in both scenarios, enabling it to handle more irregular local navigation situations. In some instances, Extreme Parkour deviates from the commands and approaches obstacles that could have been avoided. Heatmaps in Fig. 5 illustrate the position visiting frequencies of different methods in the omni-traverse task. Both Ours-s1 and Extreme Parkour struggle to navigate irregular obstacles, whereas our method visits farther positions more frequently, demonstrating more powerful locomotion skills. In the omni-traverse task with complex terrains and high obstacles, although our method may encounter collisions at terrain edges, leading to a certain probability of failure, it can still successfully accomplish the task in most cases.
@@ 4.4 Navigation System
Fig. 6 showcases sequential snapshots of our system navigating in a simulation using GPT-4[67], and in a real-world scenario using the A* algorithm. In the simulation, the navigation terrain consists of multiple terrain units, each 2m 2m, featuring varied heights and obstacles. The robot is initialized in one terrain unit, with another designated as the navigation target. The LLM (specifically, GPT-4) receives prompts as described in 3.3.2 and outputs waypoints in the form of terrain unit indices. These waypoints, positioned at the centers of the corresponding terrain units, guide our low-level controller to reach the target while handle obstacles not accounted for by the LLM. However, this waypoint configuration often overlooks intricate terrain details, potentially leading to impractical waypoint placements, such as within high obstacles. Our low-level controller can struggle to recover from some of these bad cases such as a waypoint in the middle of the gap or at the edge of a box, even though it has not been trained for such anomalies. Specifically, if the waypoint to be tracked by the robot is located in a gap region, the robot may experience partial leg suspension due to the anomalous waypoint. To maintain balance, the robot will either straddle the center of the gap or position itself close to the gap edge, rather than falling directly into the gap and causing termination.
@@ 5 Conclusion
In this work, we introduced the use of waypoints as an interface to connect a versatile quadrupedal robot with a high-level planner for navigation. The locomotion policy trained in two simulated scenarios is able to traverse complex terrains while avoiding obstacles, and the high-level planner such as LLMs enhances generalization across various environments. Experiments demonstrate that the proposed navigation architecture is able to complete navigation tasks in complex scenarios. Future works include designing an edge collision-free low-level controller for more robust locomotion and developing an end-to-end policy to conduct locomotion and navigation simultaneously.
Abbreviations
| LLM | Large language models |
| MPC | Model predictive control |
| RL | Reinforcement learning |
| RNN | Recurrent neural network |
| PPO | Proximal policy optimization |
| MDP | Markov decision process |
| WP-Fixed | Fixed waypoints |
| WP-Random | Random waypoints |
| PD | Proportional-Derivative |
| SR | Success rate |
| ATD | Average travel distance |
| AST | Average success time |
Code availability
The codes of baseline methods are from the open-source works. The codes of Skill-Nav are available from the first author.